OpenStudy (anonymous):

Which of the following is equivalent to the proportion 3/4 = x/12 ?

OpenStudy (anonymous):

boo. i can't remember this one. 4/3 = x/12 ? =/

OpenStudy (compassionate):

\(\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\) Hello, my name is Stephen Kent! I will be your tutor. /\[\frac{ 4 }{ 3 } = \frac{ x }{ 12 }\] The first step to solving the equation is to cross-multiply all terms. Can you do that for me?

OpenStudy (anonymous):

so is that 4*12 then 3*x ?

OpenStudy (compassionate):

Correct! Can you write out the new equation?

OpenStudy (anonymous):

isn't that the equation though?

OpenStudy (compassionate):

Now. The new equation is: \[48 = 3x\]

OpenStudy (anonymous):

ohh. okay. so then x = 16 ?

OpenStudy (compassionate):

Yup, that's correct!

OpenStudy (anonymous):

so then how do i get the answer for this?